OpenStudy (anonymous):

Show that 4ax ^{3}+3bx ^{2}+2cx=a+b+c has at least one root between 0 and 1.

myininaya (myininaya):

\[f(x)=4ax^3+3bx^2+2cx-a-b-c\]

OpenStudy (anonymous):

try \[f(0)\] and \[f(1)\]

myininaya (myininaya):

f(0)=-a-b-c=-(a+b+c) f(1)=4a+3b+2c-a-b-c=3a+2b+c

myininaya (myininaya):

are a, b,and c positive?

myininaya (myininaya):

if so we are done

OpenStudy (anonymous):

oh actually this is not that straightforward is it?

myininaya (myininaya):

because a+b+c is positive and -(a+b+c) is negative and 3a+2b+c is positive you know if a,b,c are all positive

OpenStudy (anonymous):

but.. suppose not. can it be the case that for any a, b, c if \[-(a+b+c)\] has one sign then \[3a+2b+c\] has opposite sign?

myininaya (myininaya):

maybe

myininaya (myininaya):

lets see if i can find a counterexample

myininaya (myininaya):

and i'm going to assume they can't all be zero

OpenStudy (anonymous):

now i am wondering if this is a set up for mvt somehow...

myininaya (myininaya):

any thoughts kwenisha?

myininaya (myininaya):

i think i found a counterexample satellite

myininaya (myininaya):

oh wait not yet lol

OpenStudy (anonymous):

after a beer i cannot, but you have three variables so it seems likely that it is possible

OpenStudy (anonymous):

i mean find three numbers a, b, c where \[a+b+c>0\] and \[3a+2b+c<0\]

OpenStudy (anonymous):

how about a = -10 b = 9, c = 2

OpenStudy (anonymous):

then \[a+b+c>0\] and \[3a+2b+c<0\] so both \[-(a+b+c)<0\] and also \[3a+2b+c<0\]

OpenStudy (anonymous):

ok maybe we can try it with the mean value theorem

OpenStudy (anonymous):

let \[F(x)=ax^4+bx^3+cx^2\] now \[F(1)=a+b+c\] \[F(0)=0\] to there exists a number "r" in [0,1] with \[F'(r)=\frac{F(1)-F(0)}{1-0}=a+b+c\] and that gives it to us because \[\frac{F(1)-F(0)}{1-0}=a+b+c\] and \[F'(x)=f(x)=4ax^3+3bx^2+2cx\]

OpenStudy (anonymous):

so mean value theorem what the right way to go, and we know there is a number "r" in [0,1] that satisfies that equation